OpenStudy (anonymous):

Given the function g(x) = 8x − 2, find g(−2).

Parth (parthkohli):

Just substitute the value of x with -2.

OpenStudy (anonymous):

All "x" with -2?

OpenStudy (anonymous):

g(-2) = 8(-2)-2 =-16-2 =-18

Parth (parthkohli):

Yes.

Parth (parthkohli):

@kevsturge yes. correct. I guess Meghan was supposed to solve this though

OpenStudy (anonymous):

Okay, Thanks again! (: I think that this was definitely my last question. Everyone here is so helpful! Thanks!

OpenStudy (anonymous):

just trying to help :D

OpenStudy (anonymous):

sometimes learning is faster if you see the answer and work backwards

OpenStudy (anonymous):

I had gotten the answer as well after the equation was written down. I just get so confused!

OpenStudy (anonymous):

thats why an estimate helps, or do the question twice and see if you get the same answer. If not then find where you did something different :D